# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 3.1 Preliminary remarks
Recall that the count of real plane curves of a given degree and genus becomes non-invariant as long as the genus is positive [11, Theorem 3.1]. However, in the case of real rational surfaces with a disconnected real part, Welschinger invariants survive for certain positive genera $g$, provided that some conditions on constraints are fulfilled: in particular, each real connected component of the ambient surface contains at most one oval (i.e., global real branch) of any curve under consideration, and the counted curves have the maximal possible number $g+1$ of ovals; see [23]. We exploit here a similar approach in order to extend the invariant count of real rational curves [19] to positive genera. Our idea is to consider real curves that are tangent to toric divisors with even multiplicity at each intersection point. In such a case, every quadrant of a real toric surface plays the role of a separate component of the real part of a real surface. The lack of the “bad” wall in the space of constraints indicated in [11, Section 3] is expressed as a cohomology vanishing condition which we prove below in Lemmas 4.1, 5.1, and 6.1.
@@ 3.2 Refined rational invariants
$g=0,\quad n_{in}=0,\quad n_{\partial}=n=\sum_{\sigma\in P^{1}_{\Delta}}n^{\sigma}-1,\quad\text{and}\quad{\boldsymbol{w}}={\boldsymbol{w}}_{\partial}\subset\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta}).$
In particular, for all $\sigma\in P^{1}_{\Delta}$ but one, $n(\sigma)=n^{\sigma}$, and for the remaining edge $\tau$, we have $n(\tau)=n^{\tau}-1$. Note that for a given ${\boldsymbol{w}}$, there exists a unique point $w^{\tau}_{n^{\tau}}\in\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ such that the sequence $\widehat{\boldsymbol{w}}:={\boldsymbol{w}}\cup\{w^{\tau}_{n^{\tau}}\}$ belongs to the $\operatorname{Men}(\Delta)$. We denote by ${\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})\subset{\mathcal{M}}_{0,n}(\Delta,{\boldsymbol{w}})$ the subset formed by the curves passing through $\widehat{\boldsymbol{w}}$.
Denote by $\operatorname{Men}^{\prime}_{0}(\Delta)$ the subset of $\operatorname{Men}(\Delta)$ formed by the above sequences $\widehat{\boldsymbol{w}}$ such that ${\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})}$
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)$, and for the sake of notation, we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$. At last, we put
$\operatorname{Men}_{0}(\Delta)=\left\{\widehat{\boldsymbol{w}}\in\operatorname{Men}^{\prime}_{0}(\Delta)\ :\ \overline{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})\right\}.$
@@
Furthermore, if $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)$, then (see [13, Section 1.1]), there is a well-defined Welschinger sign
@@
where $e_{+}(C)$ is the number of elliptic nodes of $C={\Phi}({\mathbb{P}}^{1})$ in the positive quadrant ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$ that arise from a real nodal equigeneric deformation of all singular points of $C$ in ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$, and ${\varepsilon}(\xi,w_{i}^{\sigma})$ equals $1$ or $-1$ according as the complex orientation of ${\mathbb{R}}C={\Phi}({\mathbb{P}}_{\mathbb{R}}^{1})$ at $w_{i}^{\sigma}$ agrees or not with the fixed orientation of $\partial{\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$.
@@
$W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{0}(\xi)\ .$
@@
For each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the value $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ does not depend on the choice of a generic sequence $\widehat{\boldsymbol{w}}\subset\operatorname{Men}(\Delta)$.
We prove this theorem in Section 4.
@@
Theorem 3.2 slightly generalizes the statement of [19, Theorem 5].
Remark 3.4
One can show ([19, 16]) that the invariants $W_{0}^{\kappa}(\Delta)$ vanish for all $\kappa\not\in 2{\mathbb{Z}}$ and for all $\kappa\in 2{\mathbb{Z}}$ such that $\kappa\not\equiv{\mathcal{A}}(\Delta)\mod 4$. In general, the invariants $W_{0}^{\kappa}(\Delta)$ can be computed via the formula in [2, Theorem 3.4].
@@ 3.3 Refined elliptic invariants
where ${\mathcal{Q}}$ is an open quadrant in ${\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}\setminus{\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$. As in the preceding section, for all $\sigma\in P^{1}_{\Delta}$ but one, $n(\sigma)=n^{\sigma}$, and for the remaining edge $\tau$, we have $n(\tau)=n^{\tau}-1$. As in Section 3.2, we define the completion $\widehat{\boldsymbol{w}}_{\partial}:={\boldsymbol{w}}_{\partial}\cup\{w^{\tau}_{n^{\tau}}\}\in\operatorname{Men}(\Delta)$, where $w^{\tau}_{n^{\tau}}\in\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$, and put $\widehat{\boldsymbol{w}}=\widehat{\boldsymbol{w}}_{\partial}\cup{\boldsymbol{w}}_{in}$. We denote by ${\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})\subset{\mathcal{M}}_{1,n}(\Delta,{\boldsymbol{w}})$ the subset formed by the curves passing through $\widehat{\boldsymbol{w}}$.
Denote by $\operatorname{Men}^{\prime}_{1}(\Delta)$ the set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times{\mathcal{Q}}$ such that the set ${\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})}$
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\subset\operatorname{Men}(\Delta)\times{\mathcal{Q}}$, and we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$. At last, we introduce
$\operatorname{Men}_{1}(\Delta)=\left\{\widehat{\boldsymbol{w}}\in\operatorname{Men}^{\prime}_{1}(\Delta)\ :\ \overline{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})\right\}.$
@@
$[{\Phi}:({\mathbb{E}},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})],\;\;\;\text{where}\;{\mathbb{E}}\;\text{is a smooth elliptic curve},$
defining real curves $C={\Phi}({\mathbb{E}})\subset{\operatorname{Tor}}(P_{\Delta})$ that have a one-dimensional real branch in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and that intersect the toric divisor ${\operatorname{Tor}}(\sigma)$ in $w^{\sigma}_{i}$, $1\leq i\leq n^{\sigma}$, for all $\sigma\in P^{1}_{\Delta}$. It follows from Lemma 2.4 that for $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$, the curve $C$ has two one-dimensional branches, one in ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$ and the other in the closed quadrant containing the point $w_{0}$. This defines a unique real structure on ${\mathbb{E}}$ whose fixed point set ${\mathbb{R}}{\mathbb{E}}$ consists of two ovals, making ${\Phi}:{\mathcal{E}}\to C$ a separating real curve. The choice of a component of ${\mathbb{E}}\setminus{\mathbb{R}}{\mathbb{E}}$ defines a complex orientation on the one-dimensional branches of ${\mathbb{R}}C$.
We impose a further restriction to the choice of the quadrant ${\mathcal{Q}}\supset\{w_{0}\}$ dictated by our wish to avoid specific wall-crossing events that may occur in the proof of the invariance of the count to be specified below: first, the escape of a real one-dimensional branch of counted curves out of the positive quadrant and, second, degenerations in which the union of the toric divisors splits off.
@@
Let $d$, $d_{1}$ and $d_{2}$ be positive integers. Condition (AQC) imposes no restriction on the choice of ${\mathcal{Q}}$ in the case of the triangle ${\operatorname{conv}}\{(0,0),(d,0),(0,d)\}$ (the convex hull of the points $(0,0)$, $(d,0)$, $(0,d)$) having the projective plane ${\mathbb{P}}^{2}$ as associated toric surface, while for the rectangle ${\operatorname{conv}}\{(0,0),(d_{1},0),(d_{1},d_{2}),(0,d_{2})\}$ which has ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ as associated toric surface, the condition authorizes only the quadrant ${\mathcal{Q}}$ defined by the inequalities $x_{1}<0$ and $x_{2}<0$.
@@
$([{\Phi}:({\mathbb{E}},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})],{\mathbb{E}}_{+})$
such that $[{\Phi}:({\mathbb{E}},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ and ${\mathbb{E}}_{+}$ is a half of ${\mathbb{E}}\setminus{\mathbb{R}}{\mathbb{E}}$. According to [19, Theorem 1], each element $\xi$ of $\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ has a quantum index ${\operatorname{QI}}(\xi)\in\frac{1}{2}{\mathbb{Z}}$, $|{\operatorname{QI}}(\xi)|\leq{\mathcal{A}}(\Delta)$ (defined by formula (7) with the integration domain ${\mathbb{E}}_{+}$), and for ${\boldsymbol{w}}$ such that $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$, according to the argument in [13, Section 1.1], there is a well-defined Welschinger sign
@@
where $h(C,{\mathcal{Q}})$ is the number of hyperbolic nodes of $C={\Phi}({\mathbb{E}})$ in the quadrant ${\mathcal{Q}}$ that arise in any real nodal equigeneric deformation of all singularities of $C$ in ${\mathcal{Q}}$ (here $e_{0}(C)$ and ${\varepsilon}(\xi,w_{i}^{\sigma})$ are straightforward analogs of the corresponding ingredients in formula (8)).
@@
$W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{1}(\xi)\ .$ (10)
@@
The value $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ does not depend on the choice of a generic $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times{\mathcal{Q}}$ subject to the following conditions: $\widehat{\boldsymbol{w}}_{\partial}\subset\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and $w_{0}\in{\mathcal{Q}}$, where ${\mathcal{Q}}$ is a fixed non-positive quadrant satisfying (AQC).
We prove this theorem in Section 5.
@@
One can show [16] that the invariants $W_{1}^{\kappa}(\Delta,(\alpha,\beta))$ vanish for all $\kappa\not\in 2{\mathbb{Z}}$ and for all $\kappa\in 2{\mathbb{Z}}$ such that $\kappa\not\equiv{\mathcal{A}}(\Delta)\mod 4$.

### v2

@@ 3.1 Preliminary remarks
Recall that the Welschinger signed count of real plane curves of a given degree and genus, passing through a generic conjugation-invariant configuration of points in the plane, becomes non-invariant as long as the genus is positive [14, Theorem 3.1]. However, in the case of real rational surfaces with a disconnected real part, Welschinger invariants survive for certain positive genera $g$, provided that some conditions on constraints are fulfilled: in particular, each real connected component of the ambient surface contains at most one oval (i.e., global real branch) of any curve under consideration, and the counted curves have the maximal possible number $g+1$ of ovals; see [29]. We exploit here a similar approach in order to extend the invariant count of real rational curves [24] to positive genera. Our idea is to consider real curves that are tangent to toric divisors with even intersection multiplicity at each intersection point. In such a case, every quadrant of a real toric surface plays the role of a separate component of the real part of a real surface. The lack of the “bad” wall in the space of constraints indicated in [14, Section 3] is expressed as a cohomology vanishing condition which we prove below in Lemmas 4.1, 5.1, and 6.1.
@@ 3.2 Refined rational invariants
$g=0,\quad n_{in}=0,\quad n_{\partial}=n=\sum_{\sigma\in P^{1}_{\Delta}}n^{\sigma},\quad\text{and}\quad{\boldsymbol{w}}={\boldsymbol{w}}_{\partial}\subset\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta}).$
In particular, for all $\sigma\in P^{1}_{\Delta}$ but one, we have $n(\sigma)=n^{\sigma}$, and for the remaining edge $\sigma^{\prime}$, we have $n(\sigma^{\prime})=n^{\sigma^{\prime}}-1$. Note that for a given ${\boldsymbol{w}}$, there exists a unique point $w^{\sigma^{\prime}}_{n^{\sigma^{\prime}}}\in\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ such that the sequence $\widehat{\boldsymbol{w}}:={\boldsymbol{w}}\cup\{w^{\sigma^{\prime}}_{n^{\sigma^{\prime}}}\}$ belongs to the hypersurface $\operatorname{Men}(\Delta)$. We denote by ${\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})\subset{\mathcal{M}}_{0,n}(\Delta,{\boldsymbol{w}})$ the subset formed by the curves passing through $\widehat{\boldsymbol{w}}$.
Denote by $\operatorname{Men}_{0}(\Delta)$ the subset of $\operatorname{Men}(\Delta)$ formed by the above sequences $\widehat{\boldsymbol{w}}$ such that ${\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})}\subset\overline{\mathcal{M}}^{\mathbb{R}}_{0,n}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)$, and for the sake of notation, we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$.
@@
Furthermore, if $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)$, then there are well-defined Welschinger signs
@@
[…]
where $e_{+}(C)$, respectively, $e(C)$, is the number of elliptic nodes (i.e., real intersection points of a pair of smooth transversal complex conjugate local branches) of $C={\Phi}({\mathbb{P}}^{1})$ that arise from a real nodal equigeneric deformation of all singular points of $C$ in the positive quadrant ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$, respectively, in the torus ${\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}$ (cf. [16, Section 1.1]), and ${\operatorname{sign}}(\xi,w_{i}^{\sigma})$ equals $1$ or $-1$ according as the complex orientation of ${\mathbb{R}}C={\Phi}({\mathbb{P}}_{\mathbb{R}}^{1})$ at $w_{i}^{\sigma}$ agrees or not with the fixed orientation of $\partial{\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$.
@@
$W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{0}(\xi),\quad\widetilde{W}_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}\widetilde{W}_{0}(\xi)\ .$
@@
For each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the values $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ and $\widetilde{W}_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ do not depend on the choice of a sequence $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)$.
We prove this theorem in Section 4. Since the values $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ and $\widetilde{W}_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ do not depend on the choice of a sequence $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{0}(\Delta)$, we use for these values the notations $W_{0}^{\kappa}(\Delta)$ and $\widetilde{W}_{0}^{\kappa}(\Delta)$, respectively.
@@
(1) Theorem 3.2 slightly generalizes the statement of [24, Theorem 5], because we allow tangencies of any even intersection multiplicity.
(2) One can show ([24, 19]) that the invariants $W_{0}^{\kappa}(\Delta)$ and $\widetilde{W}^{\kappa}_{0}(\Delta)$ vanish for all $\kappa\not\in 2{\mathbb{Z}}$ and for all $\kappa\in 2{\mathbb{Z}}$ such that $\kappa\not\equiv{\mathcal{A}}(\Delta)\mod 4$. In general, the invariants $W_{0}^{\kappa}(\Delta)$ can be computed via the formula in [2, Theorem 3.4].
@@ 3.3 Refined elliptic invariants
where ${\mathcal{Q}}$ is an open quadrant in ${\operatorname{Tor}}_{\mathbb{R}}(P_{\Delta})^{\times}\setminus{\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$. As in the preceding section, for all $\sigma\in P^{1}_{\Delta}$ but one, we have $n(\sigma)=n^{\sigma}$, and for the remaining edge $\sigma^{\prime}$, we have $n(\sigma^{\prime})=n^{\sigma^{\prime}}-1$. As in Section 3.2, we define the completion $\widehat{\boldsymbol{w}}_{\partial}:={\boldsymbol{w}}_{\partial}\cup\{w^{\sigma^{\prime}}_{n^{\sigma^{\prime}}}\}\in\operatorname{Men}(\Delta)$, where $w^{\sigma^{\prime}}_{n^{\sigma^{\prime}}}\in\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$, and put $\widehat{\boldsymbol{w}}=\widehat{\boldsymbol{w}}_{\partial}\cup{\boldsymbol{w}}_{in}$. We denote by ${\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})\subset{\mathcal{M}}_{1,n}(\Delta,{\boldsymbol{w}})$ the subset formed by the curves passing through $\widehat{\boldsymbol{w}}$.
Denote by $\operatorname{Men}_{1}(\Delta)$ the set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times{\mathcal{Q}}$ such that the set ${\mathcal{M}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})}\subset\overline{\mathcal{M}}^{\mathbb{R}}_{1,n}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\subset\operatorname{Men}(\Delta)\times{\mathcal{Q}}$, and we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{1,n}(\Delta,\widehat{\boldsymbol{w}})$.
@@
$[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})],\;\;\;\text{where}\;\widehat{C}\;\text{is a smooth elliptic curve},$
defining real curves $C={\Phi}(\widehat{C})\subset{\operatorname{Tor}}(P_{\Delta})$ that have a one-dimensional real branch in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and that intersect the toric divisor ${\operatorname{Tor}}(\sigma)$ in $w^{\sigma}_{i}$, $1\leq i\leq n^{\sigma}$, for all $\sigma\in P^{1}_{\Delta}$. It follows from Lemma 2.4 that for $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$, the curve $C$ has two one-dimensional branches, one in ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$ and the other in the closed quadrant $\overline{\mathcal{Q}}$ containing the point $w_{0}$. This defines a unique real structure on $\widehat{C}$ whose fixed point set ${\mathbb{R}}\widehat{C}$ consists of two ovals, making ${\Phi}:\widehat{C}\to C$ a separating real curve. The choice of a component of $\widehat{C}\setminus{\mathbb{R}}{\widehat{C}}$ defines a complex orientation on the one-dimensional branches of ${\mathbb{R}}C$.
We impose a further restriction on the choice of the quadrant ${\mathcal{Q}}\ni w_{0}$ dictated by our wish to avoid specific wall-crossing events that may occur in the proof of the invariance of the count to be specified below: first, the escape of a real one-dimensional branch of counted curves out of the positive quadrant and, second, degenerations in which the union of the toric divisors splits off.
@@
Let $d$, $d_{1}$ and $d_{2}$ be positive integers. Condition (AQC) imposes no restriction on the choice of ${\mathcal{Q}}$ in the case of the triangle ${\operatorname{conv}}\{(0,0),(d,0),(0,d)\}$ (the convex hull of the points $(0,0)$, $(d,0)$, $(0,d)$) having the projective plane ${\mathbb{P}}^{2}$ as associated toric surface, while for the rectangle ${\operatorname{conv}}\{(0,0),(d_{1},0),(d_{1},d_{2}),(0,d_{2})\}$ which has ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ as associated toric surface, the condition authorizes only the quadrant ${\mathcal{Q}}$, where the both coordinates are negative.
@@
$([{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})],\widehat{C}_{+})$
such that $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ and $\widehat{C}_{+}$ is a half of $\widehat{C}\setminus{\mathbb{R}}\widehat{C}$. According to [24, Theorem 1], each element $\xi$ of $\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})$ has a quantum index ${\operatorname{QI}}(\xi)\in\frac{1}{2}{\mathbb{Z}}$, $|{\operatorname{QI}}(\xi)|\leq{\mathcal{A}}(\Delta)$ (defined by formula (8) with the integration domain $\widehat{C}_{+}$), and for ${\boldsymbol{w}}$ such that $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)$, according to the argument in [16, Section 1.1], there is a well-defined Welschinger sign
@@
where $h(C,{\mathcal{Q}})$ is the number of hyperbolic nodes of $C={\Phi}(\widehat{C})$ in the quadrant ${\mathcal{Q}}$ that arise in any real nodal equigeneric deformation of all singularities of $C$ in ${\mathcal{Q}}$ (here $e_{0}(C)$ and ${\operatorname{sign}}(\xi,w_{i}^{\sigma})$ are straightforward analogs of the corresponding ingredients in formula (9)).
@@
$W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{1}(\xi),\quad\widetilde{W}_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}\widetilde{W}_{1}(\xi)\ ,$ (12)
where (cf. formula (10))
[…]
@@
Suppose that $n\geq 4$. Then, the values $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ and $\widetilde{W}^{\kappa}_{1}(\Delta,\widehat{\boldsymbol{w}})$ do not depend on the choice of a sequence $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)\times{\mathcal{Q}}$ subject to the following conditions: $\widehat{\boldsymbol{w}}_{\partial}\subset\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and $w_{0}\in{\mathcal{Q}}$, where ${\mathcal{Q}}$ is a fixed non-positive quadrant satisfying (AQC).
We prove this theorem in Section 5. Since the values $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ and $\widetilde{W}_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ do not depend on the choice of a sequence $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{1}(\Delta)\times{\mathcal{Q}}$, we use for these values the notations $W_{1}^{\kappa}(\Delta,{\mathcal{Q}})$ and $\widetilde{W}_{1}^{\kappa}(\Delta,{\mathcal{Q}})$, respectively.
@@
(1) We expect that the statement of Theorem 3.6 holds for the remaining value $n=3$ too. It corresponds to #$\#\Delta=3$ and $P_{\Delta}$ a lattice triangle, while the counted elliptic curves in ${\operatorname{Tor}}(P_{\Delta})$ intersect each toric divisor at one point, where they are unibranch.
(2) One can show [19] that the invariants $W_{1}^{\kappa}(\Delta,{\mathcal{Q}})$ and $\widetilde{W}_{1}^{\kappa}(\Delta,{\mathcal{Q}})$ vanish for all $\kappa\not\in 2{\mathbb{Z}}$ and for all $\kappa\in 2{\mathbb{Z}}$ such that $\kappa\not\equiv{\mathcal{A}}(\Delta)\mod 4$.